OpenStudy (anonymous):

A ball accelerates uniformly at -2.00 m/s^2. If the ball was initially moving at 0.20 m/s and the ball travels for 7.00 s, then what was its final velocity? A. -1.8 m/s B. -14.2 m/s C. +13.8 m/s D. -13.8 m/s How do I calculate this? @perl :)

OpenStudy (perl):

we can use the formula change in velocity = average acceleration * time or vf - vi= a*t or vf = a*t + vi

OpenStudy (anonymous):

okay, so is average acceleration -2.00 m/s^2 and the time is 7.00 s ?

OpenStudy (perl):

vf = (-2)*7 + 0.2

OpenStudy (anonymous):

ahh okay, thanks so much!! :D